OpenStudy (anonymous):

Solve. Express irrational solutions in simplest radical form. If the equations has no real-number solutions, write "no real-number solution." 1) (y-7)^2 = 0 2) a^2 = -9 3) 6t^2 = 156

OpenStudy (anonymous):

1) \(y=7\) 2) No real solution 3) \(6t^2=156 \implies t^2=26 \implies t=\sqrt{26}\) ot \(t=-\sqrt{26}\).

OpenStudy (anonymous):

or*